# Towards a universal QAOA protocol: Evidence of quantum advantage in solving combinatorial optimization problems

### v1

@@ Abstract
The quantum approximate optimization algorithm (QAOA) is a promising algorithm for solving combinatorial optimization problems (COPs). In this algorithm, there are alternating layers consisting of a mixer and a problem Hamiltonian. Each layer $i=0,\ldots,p-1$ is parameterized by $\beta_{i}$ and $\gamma_{i}$. How to find these parameters has been an open question with the majority of the research focused on finding them using classical algorithms. In this work, we present evidence that fixed linear ramp schedules constitute a universal set of QAOA parameters, i.e., a set of $\gamma$ and $\beta$ parameters that rapidly approximate the optimal solution, $x^{*}$, independently of the COP selected, and that the success probability to find it, $probability(x^{*})$, increases with the number of QAOA layers $p$. We simulate linear ramp QAOA protocols (LR-QAOA) involving up to $N_{q}=42$ qubits and $p=400$ layers on random instances of 9 different COPs. The results suggest that $probability(x^{*})\approx 1/2^{(\eta N_{q}/p)}$ for a constant $\eta$. We extend the analysis in 4 COPs with $p=N_{q}$ and show that $probability(x^{*})$ seems to be constant for general cases. For example, when implementing LR-QAOA with $p=42$, the $probability(x^{*})$ for 42-qubit Weighted MaxCut problems (W-MaxCut) increases from $2/2^{42}\approx 10^{-13}$ to an average of 0.13. We compare LR-QAOA, simulated annealing (SA), and branch-and-bound (B&B) finding a fundamental improvement in LR-QAOA. We test LR-QAOA on real hardware using IonQ Aria, Quantinuum H2-1, IBM Brisbane, IBM Kyoto, and IBM Osaka, encoding random weighted MaxCut (W-MaxCut) problems from 5 to 109 qubits and $p=3$ to $100$. We find that even for the largest case, $N_{q}=109$ qubits and $p=100$, information about the LR-QAOA optimization protocol is still present. The circuit involved requires 21200 CNOT gates and a time of $\approx 132\ \mu s$. The resilience of LR-QAOA to noise is attributed to its characteristic of bringing the system towards a minimal energy state despite noise driving it towards a maximally mixed state. These results show that LR-QAOA is effective in finding high-quality solutions for COPs and suggest an advantage of quantum computation for combinatorial optimization in the near future.
@@
Combinatorial Optimization, QAOA, LR-QAOA, traveling salesman, maximal independent set, bin packing, portfolio optimization, weighted maxcut, knapsack, MAX-3-SAT, .
@@ I Introduction
In this paper, we extend the study of LR-QAOA schedules to different COPs presenting analytical, numerical, and experimental evidence that LR-QAOA constitutes a universal QAOA protocol, i.e., the set of $\gamma$ and $\beta$ parameters from a linear ramp schedule that work effectively independently of the problem or problem size in combinatorial optimization. We test this protocol using maximal independent set (MIS), bin packing (BPP), traveling salesman (TSP), maximum cut (MaxCut), weighted maximum cut (W-MaxCut), 3-regular graph maximum cut (3-MaxCut), Knapsack (KP), portfolio optimization (PO), 2 Boolean satisfiability problem (Max-2-SAT), and maximum 3-SAT (Max-3-SAT). We use random instances of these COPs with problem sizes ranging from 4 to 42 qubits and $p$ from 3 to 400. For large problems, we simulate them using JUQCS–G software [17] on JUWELS Booster, a cluster of 3744 NVIDIA A100 Tensor Core GPUs, integrated into the modular supercomputer JUWELS [20, 21].
@@
We find a fundamental improvement in terms of the TTS when using LR-QAOA with the probability of success on average remaining constant $probability(x^{*})=1/2^{\eta}$ using $p=N_{q}$. With this further evidence, we conjecture that the expected probability of success will follow the relation $probability(x^{*})=1/2^{(\eta N_{q}/p)}$ for general cases of COPs. If this conjecture holds, it implies that there is a quantum algorithm that in polynomial time can solve instances of NP-Hard problems that are already hard for classical algorithms. However, we note that for some cases, the relation $probability(x^{*})=1/2^{(\eta N_{q}/p)}$ seems not to hold. We attributed it to the high concentration of sub-optimal solution close to the optimal. Even if the conjecture does not hold, there is a noticeable scaling advantage in LR-QAOA compared to these classical solvers.
We extend the analysis to real quantum hardware. Using IonQ Aria (ionq_aria), Quantinuum H2-1 (quantinuum_H2) [24], IBM Brisbane (ibm_brisbane), IBM Kyoto (ibm_kyoto), and IBM Osaka (ibm_osaka), we run W-MaxCut problems ranging from 5 to 109 qubits and $p$ from 3 to 100. We find that there is an effective number of layers, $p_{\mathrm{eff}}$, for which the best performance is obtained using each device. In the case of IBM devices $p_{\mathrm{eff}}=10$, on ionq_aria $p_{\mathrm{eff}}=10$, and on quantinuum_H2 $p_{\mathrm{eff}}=50$. Remarkably, for the largest problem size, 109 qubits and $p=100$, we observe some information about the LR-QAOA optimization is still preserved in ibm_kyoto and ibm_osaka. For a comparative analysis between the different vendors, we test a 25-qubit W-MaxCut problem on them, quantinuum_H2 gives the best performance with a probability($x^{*}$) $=0.08$ at $p=50$.
We demonstrate that LR-QAOA is resilient against noise. This robustness is attributed to the interplay of two opposing forces. While noise pushes the system towards a maximally mixed state, LR-QAOA drives the system towards the minimum energy of the cost Hamiltonian correcting errors on the way. To explore this phenomenon, we conduct simulations using a Qiskit Aer noise model [25] on a 10-qubit W-MaxCut problem. We vary the strengths of depolarizing noise affecting the 2-qubit gates within the algorithm. Our findings show significant alignment with the behavior exhibited by ionq_aria when solving the same problem.
@@ II.1 LR-QAOA
QAOA consists of alternating layers that encode the problem of interest along with a mixer element in charge of amplifying the desired solutions with low energy. In this case, the COP cost Hamiltonian is given by
@@
$U_{C}(H_{C},\gamma)=e^{-j\gamma_{i}H_{C}},$ (2)
where $\gamma_{i}$ is a parameter that in our case comes from the linear ramp schedule. Following this, in every second layer, a unitary operator is applied given by
$U(H_{B},\beta)=e^{j\beta_{i}H_{B}},$ (3)
@@
for $i=0,...,p-1$. For our simulations and experimental results, we use $\Delta_{\beta}=0.3$ and $\Delta_{\gamma}=0.6$. In the extended simulations for the conjecture, we use $\Delta_{\beta}=\Delta_{\gamma}=\Delta_{\gamma,\beta}=0.5$ and $\Delta_{\gamma,\beta}=0.4$. The special property of a linear ramp schedule is the constant slope $\Delta_{\gamma}/p$.
@@ II.2 Universal properties of LR-QAOA
where $E_{k}=\langle k|H_{c}|k\rangle$, and $|k\rangle$ is the state $k$, e.g., $|0010\rangle$. This evolution is shown in Fig. 2 (a). Equation 7 explains why the amplitude amplification is proportional to the energy of a given solution. Negative energies are rotated counterclockwise with the rotation proportional to their energies. This can be seen in Fig. 2-(e).
@@
$k\cdot l=\sum_{m=0}^{N_{q}-1}(k_{2}[m]+l_{2}[m])\mod 2,$ (9)
where $k_{2}$ ($l_{2}$) is the binary representation of $k$ ($l$). Equation 9 gives the Hamming distance between the state $k$ and $l$. For example, in a 3-qubit system, the evolution of $\alpha^{t}_{000}$ is given by
@@
A schematic representation of how the $U_{B}(\beta_{t})$ induces an evolution of $\alpha_{000}$ is shown in Fig. 2-(b). Here, $U_{B}(\beta_{t})$ changes the amplitude and direction of $\alpha_{000}$ using the information of this and the other states. The Hamming distance indicates how the amplitudes are grouped. For example, the effective vector $r_{1}=(\alpha_{001}+\alpha_{010}+\alpha_{100})$ of states with Hamming distance 1, contribute to $\alpha_{000}$ after a $\pi/2$ rotation and a rescaling given by $\sin(\beta_{t})\cos^{2}(\beta_{t})$.
In Fig. 2-(c) is shown the evolution of the optimal solution, $x^{*}$, of a 6-qubit W-MaxCut problem for the $U_{C}(\gamma_{t})$ and $U_{B}(\beta_{t})$ steps for $t\in\{0,...,Nq-1\}$. In Fig.2-(d) shows the same evolution but for the lowest eigenvalue. In this case, the evolution of the eigenvalue due to $U_{C}(\gamma_{t})$ goes in the opposite direction to the evolution of $U_{B}(\beta_{t})$ producing the desired effect of interference. Fig.2-(e) shows the angle of rotation of the white triangles, i.e., the rotation due to $\theta(\gamma_{t})$. The gain in the amplitude of the $\alpha_{k}^{t}$ at each time step after the unitary evolution $U_{B}(\beta_{t})$ is shown in Fig. 2-(f).
@@ II.5 Performance diagram
In Fig. 3-(b), we present the performance diagram of a 10-qubit W-MaxCut using LR-QAOA. The shaded red region in this plot illustrates the ’ridge’ as defined by [11]. The ridge region is characterized by consistently improving performance as the number of layers grows. The two points in this plot show similar performance, with the black point achieving a $probability(x^{*})>0.65$ with just $p=20$, while the white point in the annealing region requires a higher $p=300$.
@@ III.1 Simulations
Figure 4-(g) shows a 12-qubit PO probability vs. the cost of the first 3000 sorted eigenvalues using LR-QAOA from 10 to 200 layers. The dots represent the probability of getting a certain cost Hamiltonian eigenvalue vs. the cost associated with it. Only the first 3000 eigenvalues with the lowest cost are presented. This plot shows another characteristic of LR-QAOA, namely that the probability of obtaining a given solution drops exponentially with increasing cost. Even if the optimal solution is not observed in one sample of LR-QAOA, it is more likely that a low-cost energy is observed.
@@ III.2 Conjecture
Figure 5-(c) shows the $TTS_{99\%}$ for a $p_{d}=0.99$ in Eq. 12 of the W-MaxCut problems for SA and LR-QAOA. We include the CPLEX number of iterations needed to find the optimal solution for comparison. This plot shows the exponential scaling of classical solvers to find the optimal solution to NP-hard problems. On average, it grows with $N_{iterations}=2^{0.38N_{q}+1.5}$ for CPLEX after 20 qubits. In the case of LR-QAOA, it stabilizes at around $10^{3}$ iterations, and at $N_{q}=42$, it has improved by two orders of magnitude the TTS compared to SA. LR-QAOA has the fundamental advantage of knowing the full energy spectrum of the COP at each time step while the other methods need some time to learn it.
@@
Figure 5: Evidence for conjecutre 1. The problems shown here are solved using LR-QAOA with $p=N_{q}$ (a) Probability of success of fully connected W-MaxCut problem using 100 random W-MaxCut problems up to 40 qubits and 10 problems for 42 qubits. The different vertically stacked dots represent individual cases and the circles (triangles) indicate the mean value $\Delta_{\beta,\gamma}=0.5$ ($\Delta_{\beta,\gamma}=0.4$). The dotted line represents the probability of finding the optimal solution by random guessing and the dashed line represents the probability of finding the optimal solution if the search space is reduced by a quadratic speedup algorithm. (b) Approximation ratio for the W-MaxCut problems. The dashed line (GW) represents the approximation ratio using the best-known polynomial-time classical algorithm for the same problem. (c) TTS for SA and LR-QAOA compared to CPLEX’s number of iterations. The dashed line is added as a guiding line it represents a quadratic speedup $2^{N_{q}/2}$. (d) 20 hard cases of Max-2-SAT problems (triangles) for problem sizes up to 20 qubits and 10 random cases of Max-3-SAT problems (circles) for problems up to 25 qubits. The inset shows the degeneracies of the 4 first eigenvalues for the Max-2-SAT problem highlighted in red. (e) 10 random MaxCut instances for problem sizes up to 35 qubits. The different colors mean the average percentage of edges compared to a fully connected graph. Therefore, darker blue, $0.05$, means a $5\%$ of having edges. The circles represent the mean value and the dots represent the worst case. The inset shows the degeneracies of the solutions in the worst case. (f) The number of iterations needed to find the optimal solution using CPLEX for the MaxCut problems. The colors follow the same pattern as in (e).

### v2

# Towards a universal QAOA protocol: Evidence of a scaling advantage in solving some combinatorial optimization problems

@@ Abstract
The quantum approximate optimization algorithm (QAOA) is a promising algorithm for solving combinatorial optimization problems (COPs). In this algorithm, there are alternating layers consisting of a mixer and a problem Hamiltonian. Each layer $i=0,\ldots,p-1$ is parameterized by $\beta_{i}$ and $\gamma_{i}$. How to find these parameters has been an open question with the majority of the research focused on finding them using classical algorithms. In this work, we present evidence that fixed linear ramp schedules constitute a universal set of QAOA parameters, i.e., a set of $\gamma$ and $\beta$ parameters that rapidly approximate the optimal solution, $x^{*}$, independently of the COP selected, and that the success probability of finding it, $probability(x^{*})$, increases with the number of QAOA layers $p$. We simulate linear ramp QAOA protocols (LR-QAOA) involving up to $N_{q}=42$ qubits and $p=400$ layers on random instances of 9 different COPs. The results suggest that $probability(x^{*})\approx 1/2^{(\eta N_{q}/p)}$ for a constant $\eta$. We extend the analysis in 4 COPs with $p=N_{q}$ and show that $probability(x^{*})$ seems to be constant for general cases. For example, when implementing LR-QAOA with $p=42$, the $probability(x^{*})$ for 42-qubit Weighted MaxCut problems (W-MaxCut) increases from $2/2^{42}\approx 10^{-13}$ to an average of 0.13. We compare LR-QAOA, simulated annealing (SA), and branch-and-bound (B&B) finding a scaling improvement in LR-QAOA. We test LR-QAOA on real hardware using IonQ Aria, Quantinuum H2-1, IBM Brisbane, IBM Kyoto, and IBM Osaka, encoding random weighted MaxCut (W-MaxCut) problems from 5 to 109 qubits and $p=3$ to $100$. We find that even for the largest case, $N_{q}=109$ qubits and $p=100$, information about the LR-QAOA optimization protocol is still present. The circuit involved requires 21200 CNOT gates and a time of $\approx 132\ \mu s$. The resilience of LR-QAOA to noise is attributed to its characteristic of bringing the system towards a minimal energy state despite noise driving it towards a maximally mixed state. These results show that LR-QAOA effectively finds high-quality solutions for a large variety of COPs and suggest a scaling advantage of quantum computation for combinatorial optimization.
@@
Combinatorial Optimization, QAOA, LR-QAOA, traveling salesman, maximal independent set, bin packing, portfolio optimization, weighted MaxCut, knapsack, MAX-3-SAT.
@@ I Introduction
In this paper, we extend the study of LR-QAOA schedules to different COPs presenting numerical and experimental evidence that LR-QAOA constitutes a universal QAOA protocol, i.e., the set of $\gamma$ and $\beta$ parameters from a linear ramp schedule that work effectively independently of the problem or problem size in combinatorial optimization. We test this protocol using maximal independent set (MIS), bin packing (BPP), traveling salesman (TSP), maximum cut (MaxCut), weighted maximum cut (W-MaxCut), 3-regular graph maximum cut (3-MaxCut), Knapsack (KP), portfolio optimization (PO), maximum 2 Boolean satisfiability problem (Max-2-SAT), and maximum 3-SAT (Max-3-SAT). We use random instances of these COPs with problem sizes ranging from 4 to 42 qubits and $p$ from 3 to 400. For large problems, we simulate them using JUQCS–G software [17] on JUWELS Booster, a cluster of 3744 NVIDIA A100 Tensor Core GPUs, integrated into the modular supercomputer JUWELS [20, 21].
@@
We find a scaling improvement in terms of the TTS when using LR-QAOA with the probability of success on average remaining constant $probability(x^{*})=1/2^{\eta}$ using $p=N_{q}$. With this further evidence, we conjecture that the expected probability of success will follow the relation $probability(x^{*})=1/2^{(\eta N_{q}/p)}$ for typical cases of COPs. If this conjecture holds, it implies that there is a quantum algorithm that in polynomial time can solve instances of NP-Hard problems that are already hard for classical algorithms. This linear time complexity has also been observed recently for 3-MaxCut problems in [24] using what they called Floquet adiabatic algorithm which follows the same ansatz that LR-QAOA but from the point of view of a Trotterization approximation of the adiabatic algorithm similar to AQA. However, we note that for some cases, the relation $probability(x^{*})=1/2^{(\eta N_{q}/p)}$ seems not to hold. We attributed it to the high concentration of sub-optimal solution close to the optimal. Even if the conjecture does not hold, there is a noticeable scaling advantage in LR-QAOA compared to SA and B&B for solving instances of W-MaxCut. This evidence complements the recent findings in [25] where a scaling advantage is observed of a fixed QAOA protocol for solving a classical intractable COP known as low autocorrelation binary sequences (LABS).
We extend the analysis to real quantum hardware. Using IonQ Aria (ionq_aria), Quantinuum H2-1 (quantinuum_H2) [26], IBM Brisbane (ibm_brisbane), IBM Kyoto (ibm_kyoto), and IBM Osaka (ibm_osaka), we run W-MaxCut problems ranging from 5 to 109 qubits and $p$ from 3 to 100. We find that there is an effective number of layers, $p_{\mathrm{eff}}$, for which the best performance is obtained using each device. In the case of IBM devices $p_{\mathrm{eff}}=10$, on ionq_aria $p_{\mathrm{eff}}=10$, and on quantinuum_H2 $p_{\mathrm{eff}}=50$. Remarkably, for the largest problem size, 109 qubits and $p=100$, we observe that LR-QAOA still possesses an advantage over random sampling in ibm_kyoto and ibm_osaka. For a comparative analysis between the different vendors, we test a 25-qubit W-MaxCut problem on them, quantinuum_H2 gives the best performance with a probability($x^{*}$) $=0.08$ at $p=50$.
We demonstrate that LR-QAOA is resilient to noise. This robustness is attributed to the interplay of two opposing forces. While noise pushes the system towards a maximally mixed state, LR-QAOA drives the system towards the minimum energy of the cost Hamiltonian correcting errors on the way. To explore this phenomenon, we conduct simulations using a Qiskit Aer noise model [27] on a 10-qubit W-MaxCut problem. We vary the strengths of depolarizing noise affecting the 2-qubit gates within the algorithm. Our findings show significant alignment with the behavior exhibited by ionq_aria when solving the same problem.
@@ II.1 LR-QAOA
QAOA consists of alternating layers that encode the problem of interest along with a mixer element in charge of amplifying solutions with low energy. In this case, the COP cost Hamiltonian is given by
@@
$U_{C}(H_{C},\gamma_{i})=e^{-j\gamma_{i}H_{C}},$ (2)
where $\gamma_{i}$ is a parameter that in our case comes from the linear ramp schedule. Following this, in every second part of a layer, a unitary operator is applied given by
$U(H_{B},\beta_{i})=e^{j\beta_{i}H_{B}},$ (3)
@@
for $i=0,...,p-1$. For our simulations and experimental results, we use $\Delta_{\beta}=0.3$ and $\Delta_{\gamma}=0.6$. In the extended simulations for the conjecture, we use $\Delta_{\beta}=\Delta_{\gamma}=\Delta_{\gamma,\beta}=0.5$ and $\Delta_{\gamma,\beta}=0.4$.
@@ II.2 Universal properties of LR-QAOA
where $E_{k}=\langle k|H_{c}|k\rangle$. This evolution is shown in Fig. 2 (a). Equation 7 explains why the amplitude amplification is proportional to the energy of a given solution. Negative energies are rotated counterclockwise with the rotation proportional to their energies. This can be seen in Fig. 2-(e).
@@
$k\cdot l=\sum_{m=0}^{N_{q}-1}(k[m]+l[m])\mod 2,$ (9)
where $k$ and $l$ are states in the computational basis. Equation 9 gives the Hamming distance between the state $k$ and $l$. For example, in a 3-qubit system, the evolution of $\alpha^{t}_{000}$ is given by
@@
A schematic representation of how the $U_{B}(\beta_{t})$ induces an evolution of $\alpha_{000}$ is shown in Fig. 2-(b). Here, $U_{B}(\beta_{t})$ changes the amplitude and direction of $\alpha_{000}$ using the information of $\alpha_{000}$ and the other states. The Hamming distance indicates how the amplitudes are grouped. For example, the effective vector $r_{1}=(\alpha_{001}+\alpha_{010}+\alpha_{100})$ of states with Hamming distance 1, contribute to $\alpha_{000}$ after a $\pi/2$ rotation and a rescaling given by $\sin(\beta_{t})\cos^{2}(\beta_{t})$.
In Fig. 2-(c) is shown the evolution of the optimal solution, $x^{*}$, of a 6-qubit W-MaxCut problem for the $U_{C}(\gamma_{t})$ and $U_{B}(\beta_{t})$ steps for $t\in\{0,...,Nq-1\}$. In Fig.2-(d) shows the same evolution but for the state with the lowest energy. In this case, the evolution of the eigenvalue due to $U_{C}(\gamma_{t})$ goes in the opposite direction to the evolution of $U_{B}(\beta_{t})$ producing the desired effect of interference. Fig.2-(e) shows the angle of rotation of the white triangles, i.e., the rotation due to $\theta(\gamma_{t})$. The gain in the amplitude of the $\alpha_{k}^{t}$ at each time step after the unitary evolution $U_{B}(\beta_{t})$ is shown in Fig. 2-(f).
@@ II.5 Performance diagram
In Fig. 3-(b), we present the performance diagram of a 10-qubit W-MaxCut using LR-QAOA. The shaded red region in this plot illustrates the ’ridge’ as defined by [11]. The ridge region is characterized by consistently improving performance as the number of layers grows. The two points in this plot show similar performance, with the black point achieving a $probability(x^{*})>0.65$ with just $p=20$, while the white point in the annealing region requires a higher $p=300$. In appendix A.2, it is shown a 2-D version of the performance diagram for different problems.
@@ III.1 Simulations
Figure 4-(g) shows a 12-qubit PO probability vs. the cost of the first 3000 sorted eigenvalues using LR-QAOA from 10 to 200 layers. The dots represent the probability of getting a certain cost Hamiltonian eigenvalue vs. the cost associated with it. This plot shows another characteristic of LR-QAOA, namely that the probability of obtaining a given solution drops exponentially with increasing cost. Even if the optimal solution is not observed in one sample of LR-QAOA, it is more likely that a low-cost energy is observed.
@@ III.2 Conjecture
Figure 5-(c) shows the $TTS_{99\%}$ for a $p_{d}=0.99$ (Eq. 12) of the W-MaxCut problems for SA and LR-QAOA. We include the CPLEX number of iterations needed to find the optimal solution for comparison. This plot shows the average TTS for the 100 W-MaxCut cases. We add three guiding lines of the perceived scaling of the 3 algorithms. CPLEX B&B seems to scale $\approx 2^{0.36N_{q}}$, SA $\approx 2^{0.19N_{q}}$, and LR-QAOA $\approx 2^{0.11N_{q}}$. For LR-QAOA, we also show an exponential growth line, but based on our conjecture, it could potentially grow linearly. Note that we do not tune the $\Delta_{\gamma,\beta}$ parameters in these cases, and even though the results are promising, further improvement can be achieved finding a function $\Delta_{\gamma,\beta}(Nq)$. In the case of LR-QAOA, at $N_{q}=42$, it has improved by two orders of magnitude the TTS compared to SA and B&B.
@@
Figure 5: Evidence for conjecutre 1. The problems shown here are solved using LR-QAOA with $p=N_{q}$ (a) Probability of success of fully connected W-MaxCut problem using 100 random W-MaxCut problems up to 40 qubits and 10 problems for 42 qubits. The different vertically stacked dots represent individual cases and the circles (triangles) indicate the mean value $\Delta_{\beta,\gamma}=0.5$ ($\Delta_{\beta,\gamma}=0.4$). The dotted line represents the probability of finding the optimal solution by random guessing and the dashed line represents the probability of finding the optimal solution if the search space is reduced by a quadratic speedup algorithm. (b) Approximation ratio for the W-MaxCut problems. The dashed line (GW) represents the approximation ratio using the best-known polynomial-time classical algorithm for the same problem. (c) TTS for SA and LR-QAOA compared to CPLEX’s number of iterations. Three guiding lines are added following the perceived scale of the different algorithms. (d) 20 hard cases of Max-2-SAT problems (triangles) for problem sizes up to 20 qubits and 10 random cases of Max-3-SAT problems (circles) for problems up to 25 qubits. The inset shows the degeneracies of the 4 first eigenvalues for the Max-2-SAT problem highlighted in red. (e) 10 random MaxCut instances for problem sizes up to 35 qubits. The different colors mean the average percentage of edges compared to a fully connected graph. Therefore, darker blue, $0.05$, means a $5\%$ of having edges. The circles represent the mean value and the dots represent the worst case. The inset shows the degeneracies of the solutions in the worst case. (f) The number of iterations needed to find the optimal solution using CPLEX for the MaxCut problems. The colors follow the same pattern as in (e).